30N and 46N forces are perpendicular to each other and pull a single object what is the third force that must act on the object if the result is 0


1
Expert's answer
2021-01-24T19:06:36-0500

Let "F_1=30\ N" pulls the object along the horizontal and "F_2=46\ N" pulls the object along the vertical. Let's find the magnitude and direction of the third force. Let’s write the sum of projections of the forces on axis "x"- and "y":


"F_{1x}+F_{2x}+F_{3x}=0,""F_{1y}+F_{2y}+F_{3y}=0."

Then, we get:


"30\ N+0\ N+F_{3x}=0,""F_{3x}=-30\ N,""0\ N+46\ N+F_{3y}=0,""F_{3y}=-46\ N."

We can find the magnitude of the third force from the Pythagorean theorem:


"F_3=\sqrt{F_{3x}^2+F_{3y}^2}=\sqrt{(-30\ N)^2+(-46\ N)^2}=55\ N."

Let's find the direction of the third force:


"cos\theta=\dfrac{F_{3x}}{F_3},""\theta=cos^{-1}(\dfrac{F_{3x}}{F_3}),""\theta=cos^{-1}(\dfrac{-30\ N}{55\ N})=123^{\circ}."


In order to find correct angle we must subtruct the obtained angle from "360^{\circ}":


"\theta=360^{\circ}-123^{\circ}=237^{\circ}."

To obtaine zero resultunt force, the third force must have the magnitude 55 N and direction "237^{\circ}"(counted conterclockwise from the "x"-axis).